# To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States

### v1

@@ 4.2 Results from Ordinal Quantile Models
$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
EV Info $-0.0722$ $-0.0421$ $0.0527$ $0.0615$ $-0.0792$ $-0.0374$ $0.0172$ $0.0995$ $-0.0722$ $-0.0453$ $-0.0089$ $0.1265$
Env Better $-0.1958$ $-0.0191$ $0.1135$ $0.1014$ $-0.1771$ $-0.0730$ $0.1122$ $0.1379$ $-0.1384$ $-0.1127$ $0.0789$ $0.1722$
EV Owner $-0.1357$ $-0.0896$ $0.0633$ $0.1620$ $-0.1212$ $-0.0757$ $-0.0113$ $0.2083$ $-0.1053$ $-0.0953$ $-0.0265$ $0.2270$
GACC (TM) $0.2192$ $0.0097$ $-0.1217$ $-0.1072$ $0.1974$ $0.0635$ $-0.1203$ $-0.1406$ $0.1283$ $0.0779$ $-0.0575$ $-0.1487$
GACC (RA) $0.1216$ $0.0312$ $-0.0780$ $-0.0747$ $0.1097$ $0.0534$ $-0.0613$ $-0.1018$ $0.0763$ $0.0421$ $-0.0124$ $-0.1059$
Age $<50$ $-0.0237$ $-0.0106$ $0.0156$ $0.0187$ $-0.0303$ $-0.0134$ $0.0095$ $0.0343$ $-0.0265$ $-0.0164$ $0.0003$ $0.0427$
Metropolitan $-0.0483$ $-0.0174$ $0.0318$ $0.0338$ $-0.0565$ $-0.0225$ $0.0237$ $0.0553$ $-0.0345$ $-0.0185$ $0.0030$ $0.0500$
Midwest $0.0185$ $0.0080$ $-0.0123$ $-0.0141$ $0.0161$ $0.0068$ $-0.0053$ $-0.0177$ $0.0114$ $0.0066$ $-0.0003$ $-0.0177$
South $0.0124$ $0.0056$ $-0.0083$ $-0.0098$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
West $-0.0327$ $-0.0168$ $0.0220$ $0.0275$ $-0.0382$ $-0.0176$ $0.0092$ $0.0467$ $-0.0311$ $-0.0200$ $-0.0015$ $0.0526$
Black $0.0784$ $0.0256$ $-0.0519$ $-0.0520$ $0.0613$ $0.0244$ $-0.0259$ $-0.0598$ $0.0269$ $0.0141$ $-0.0014$ $-0.0396$
Democrat $-0.1242$ $-0.0523$ $0.0808$ $0.0958$ $-0.1081$ $-0.0631$ $0.0431$ $0.1280$ $-0.0828$ $-0.0676$ $0.0070$ $0.1434$
@@
EV Owner $-0.1782$ $-0.0723$ $0.0880$ $0.1624$ $-0.1573$ $-0.0695$ $0.0121$ $0.2147$ $-0.1465$ $-0.1110$ $-0.0071$ $0.2646$
GACC (TM) $0.3425$ $-0.0568$ $-0.1515$ $-0.1341$ $0.3304$ $0.0226$ $-0.1778$ $-0.1752$ $0.2471$ $0.0963$ $-0.1459$ $-0.1975$
GACC (RA) $0.1541$ $0.0026$ $-0.0790$ $-0.0777$ $0.1527$ $0.0497$ $-0.0852$ $-0.1171$ $0.1269$ $0.0392$ $-0.0308$ $-0.1352$
Age $<50$ $-0.0362$ $-0.0066$ $0.0199$ $0.0229$ $-0.0390$ $-0.0085$ $0.0130$ $0.0344$ $-0.0358$ $-0.0167$ $0.0072$ $0.0453$
Metropolitan $-0.0684$ $-0.0054$ $0.0360$ $0.0378$ $-0.0643$ $-0.0109$ $0.0258$ $0.0493$ $-0.0483$ $-0.0165$ $0.0114$ $0.0535$
Black $0.0642$ $0.0060$ $-0.0342$ $-0.0359$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Democrat $-0.1719$ $-0.0295$ $0.0927$ $0.1087$ $-0.1437$ $-0.0574$ $0.0601$ $0.1410$ $-0.0996$ $-0.0659$ $0.0217$ $0.1438$
@@
EV Infra $-0.1863$ $-0.0428$ $0.0956$ $0.1335$ $-0.1827$ $-0.0603$ $0.0616$ $0.1813$ $-0.1452$ $-0.0730$ $0.0187$ $0.1995$
EV Owner $-0.1684$ $-0.0350$ $0.0727$ $0.1307$ $-0.1842$ $-0.0542$ $0.0459$ $0.1925$ $-0.1580$ $-0.0817$ $0.0183$ $0.2215$
GACC (TM) $0.3323$ $-0.0635$ $-0.1405$ $-0.1283$ $0.3103$ $-0.0250$ $-0.1426$ $-0.1427$ $0.2512$ $0.0486$ $-0.1423$ $-0.1575$
GACC (RA) $0.1318$ $-0.0065$ $-0.0592$ $-0.0661$ $0.1238$ $0.0132$ $-0.0570$ $-0.0799$ $0.0975$ $0.0183$ $-0.0329$ $-0.0830$
Age $<50$ $-0.0296$ $-0.0007$ $0.0131$ $0.0173$ $-0.0363$ $-0.0034$ $0.0139$ $0.0258$ $-0.0349$ $-0.0082$ $0.0102$ $0.0329$
Metropolitan $-0.0568$ $0.0013$ $0.0250$ $0.0304$ $-0.0511$ $-0.0026$ $0.0204$ $0.0333$ $-0.0408$ $-0.0067$ $0.0129$ $0.0345$
Black $0.0665$ $-0.0010$ $-0.0296$ $-0.0359$ $0.0462$ $0.0029$ $-0.0182$ $-0.0309$ $0.0239$ $0.0039$ $-0.0069$ $-0.0209$
Democrat $-0.1832$ $-0.0084$ $0.0851$ $0.1066$ $-0.1815$ $-0.0384$ $0.0798$ $0.1401$ $-0.1393$ $-0.0660$ $0.0491$ $0.1562$
@@
Beliefs about EVs’ environmental benefits (measured in the 2021 survey) also show varying relationships with purchase intent across quantiles. Among respondents at the 80th percentile, those who believe EVs are better for the environment are 17.2 percentage points more likely to express strong purchase intent compared to those at the same percentile who do not share this belief. The corresponding differences are 13.8 percentage points at the median and 10.1 percentage points at the 20th percentile. Conversely, such respondents are 13.8 percentage points less likely to express no interest at the 80th percentile, compared to 17.7 percentage points less likely at the median and 19.6 percentage points less likely at the 20th percentile. This finding suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters while easing skepticism among consumers with lower purchase interest; this indicates that highlighting benefits such as local air quality improvements could broaden support for EV adoption.
The influence of prior EV ownership also varies across quantiles. At the 80th percentile, respondents with prior EV ownership are 22.7 percentage points more likely to express strong purchase intent compared to those without prior ownership. This difference decreases to 20.8 percentage points at the median and 16.2 percentage points at the 20th percentile. A similar pattern is observed in both the 2022 and 2023 surveys. In the 2022 survey, the probability of strong interest increases by 26.5, 21.5, and 16.2 percentage points at the 80th, 50th, and 20th quantiles, respectively. For the same quantiles, the covariate effects for strong interest in the 2023 survey are 22.1, 19.2, and 13.1 percentage points.
@@
When it comes to views on government action, respondents who believe the government is doing too much or enough to reduce climate change are significantly less likely to express strong interest in purchasing an EV, with this negative effect becoming more pronounced at higher percentiles. In the 2021 survey, those at the 80th percentile of EV purchase intent who hold this belief are 14.9 percentage points less likely to report strong purchase intent compared to those at the same percentile who do not. The gap is 14.1 percentage points at the median and 10.7 percentage points at the 20th percentile. Likewise, at the 80th percentile, these respondents are 12.8 percentage points more likely to express no interest in EVs, with the gap widening to 19.7 percentage points at the median and 21.9 percentage points at the 20th percentile. More recent surveys indicate even stronger covariate effects, with those who believe the government is doing too much on climate issues being approximately 33 percentage points more likely to express no interest in EVs at the 20th percentile and 25 percentage points more likely at the 80th percentile. These patterns suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonates across ideological lines. Rather than positioning EVs primarily as a climate solution, emphasizing other benefits—such as long-term cost savings and air quality improvements—may be more effective in engaging a broader range of potential buyers. Moreover, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns, particularly for consumers who are wary of climate-focused policies.
Figure 3: The figure presents the covariate effects for some selected variables from the April 2021 and May-June 2023 surveys, spanning the 5th to 95th quantiles with 5th quantile increments. Legend: ΔPr(Not at all $\Delta\Pr(\textrm{Not at all likely})$:  ; ΔPr(Not too $\Delta\Pr(\textrm{Not too likely})$:   ; ΔPr(Somewhat $\Delta\Pr(\textrm{Somewhat likely})$:   ; ΔPr(Very $\Delta\Pr(\textrm{Very likely})$:  .
@@
Finally, the relationship between political party affiliation and EV purchase intent varies considerably across quantiles. Results from the 2021 survey show that at the 80th percentile Democrats are 14.3 percentage points more likely than Republicans to express strong purchase intent, with the difference narrowing to 9.6 percentage points among respondents at the 20th percentile. Similarly, at the 80th percentile Democrats are 8.3 percentage points less likely than Republicans to express no interest in EVs, while the difference was 12.4 percentage points at the 20th percentile. These patterns remain consistent in more recent surveys. The results indicate that political identity is closely linked to EV purchase intent across the entire distribution: at higher quantiles, Democrats are significantly more likely to express strong purchase intent, while at lower quantiles, they are far less likely to reject EVs outright. Emphasizing benefits that resonate across political ideologies—such as long-term cost savings and energy independence—may help broaden EV appeal across the political spectrum.
@@
To provide a clearer understanding of how covariate effects vary across quantiles, Figure 3 presents the covariate effects for some selected variables from a series of ordinal quantile regression models, covering the 5th to the 95th percentiles of EV purchase intent in 5-percentile increments. Each panel corresponds to a different covariate; with the solid gray, dashed gray, dashed black, and solid black lines representing the change in probability of reporting “not likely at all,” “not too likely,” “somewhat likely,” and “very likely,” respectively, for a given change in covariate across the quantiles. The figure clearly illustrates the substantial differences in covariate effects across the distribution of EV purchase intent.
@@
For example, the solid black line in these panels indicates that respondents at higher quantiles who have read or heard extensively about EVs, perceive them as environmentally beneficial, have previously owned an EV, identify as Democrats, or are confident in the United States’ ability to develop the necessary infrastructure are much more likely to report strong interest in purchasing EVs compared to those at lower quantiles. Conversely, the solid gray lines show that those who believe government climate efforts are excessive or adequate are significantly more likely to report no interest in buying EVs at lower quantiles and markedly less likely to report strong interest in buying EVs at higher quantiles.
Overall, these results highlight the need for tailored strategies to address the barriers to EV adoption for different population segments.

### v2

@@ 4.2 Results from Ordinal Quantile Models
[…]
EV Info 0.0722-0.0722 0.0421-0.0421 $0.0527$ $0.0615$ 0.0792-0.0792 0.0374-0.0374 $0.0172$ $0.0995$ 0.0722-0.0722 0.0453-0.0453 0.0089-0.0089 $0.1265$
Env Better 0.1958-0.1958 0.0191-0.0191 $0.1135$ $0.1014$ 0.1771-0.1771 0.0730-0.0730 $0.1122$ $0.1379$ 0.1384-0.1384 0.1127-0.1127 $0.0789$ $0.1722$
GACC (TM) $0.2192$ $0.0097$ 0.1217-0.1217 0.1072-0.1072 $0.1974$ $0.0635$ 0.1203-0.1203 0.1406-0.1406 $0.1283$ $0.0779$ 0.0575-0.0575 0.1487-0.1487
GACC (RA) $0.1216$ $0.0312$ 0.0780-0.0780 0.0747-0.0747 $0.1097$ $0.0534$ 0.0613-0.0613 0.1018-0.1018 $0.0763$ $0.0421$ 0.0124-0.0124 0.1059-0.1059
EV Owner 0.1357-0.1357 0.0896-0.0896 $0.0633$ $0.1620$ 0.1212-0.1212 0.0757-0.0757 0.0113-0.0113 $0.2083$ 0.1053-0.1053 0.0953-0.0953 0.0265-0.0265 $0.2270$
Age <50<50 0.0237-0.0237 0.0106-0.0106 $0.0156$ $0.0187$ 0.0303-0.0303 0.0134-0.0134 $0.0095$ $0.0343$ 0.0265-0.0265 0.0164-0.0164 $0.0003$ $0.0427$
Metropolitan 0.0483-0.0483 0.0174-0.0174 $0.0318$ $0.0338$ 0.0565-0.0565 0.0225-0.0225 $0.0237$ $0.0553$ 0.0345-0.0345 0.0185-0.0185 $0.0030$ $0.0500$
Midwest $0.0185$ $0.0080$ 0.0123-0.0123 0.0141-0.0141 $0.0161$ $0.0068$ 0.0053-0.0053 0.0177-0.0177 $0.0114$ $0.0066$ 0.0003-0.0003 0.0177-0.0177
South $0.0124$ $0.0056$ 0.0083-0.0083 0.0098-0.0098 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
West 0.0327-0.0327 0.0168-0.0168 $0.0220$ $0.0275$ 0.0382-0.0382 0.0176-0.0176 $0.0092$ $0.0467$ 0.0311-0.0311 0.0200-0.0200 0.0015-0.0015 $0.0526$
Black $0.0784$ $0.0256$ 0.0519-0.0519 0.0520-0.0520 $0.0613$ $0.0244$ 0.0259-0.0259 0.0598-0.0598 $0.0269$ $0.0141$ 0.0014-0.0014 0.0396-0.0396
Democrat 0.1242-0.1242 0.0523-0.0523 $0.0808$ $0.0958$ 0.1081-0.1081 0.0631-0.0631 $0.0431$ $0.1280$ 0.0828-0.0828 0.0676-0.0676 $0.0070$ $0.1434$
@@
GACC (TM) $0.3425$ 0.0568-0.0568 0.1515-0.1515 0.1341-0.1341 $0.3304$ $0.0226$ 0.1778-0.1778 0.1752-0.1752 $0.2471$ $0.0963$ 0.1459-0.1459 0.1975-0.1975
GACC (RA) $0.1541$ $0.0026$ 0.0790-0.0790 0.0777-0.0777 $0.1527$ $0.0497$ 0.0852-0.0852 0.1171-0.1171 $0.1269$ $0.0392$ 0.0308-0.0308 0.1352-0.1352
EV Owner 0.1782-0.1782 0.0723-0.0723 $0.0880$ $0.1624$ 0.1573-0.1573 0.0695-0.0695 $0.0121$ $0.2147$ 0.1465-0.1465 0.1110-0.1110 0.0071-0.0071 $0.2646$
Age <50<50 0.0362-0.0362 0.0066-0.0066 $0.0199$ $0.0229$ 0.0390-0.0390 0.0085-0.0085 $0.0130$ $0.0344$ 0.0358-0.0358 0.0167-0.0167 $0.0072$ $0.0453$
Metropolitan 0.0684-0.0684 0.0054-0.0054 $0.0360$ $0.0378$ 0.0643-0.0643 0.0109-0.0109 $0.0258$ $0.0493$ 0.0483-0.0483 0.0165-0.0165 $0.0114$ $0.0535$
Black $0.0642$ $0.0060$ 0.0342-0.0342 0.0359-0.0359 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Democrat 0.1719-0.1719 0.0295-0.0295 $0.0927$ $0.1087$ 0.1437-0.1437 0.0574-0.0574 $0.0601$ $0.1410$ 0.0996-0.0996 0.0659-0.0659 $0.0217$ $0.1438$
@@
EV Infra 0.1863-0.1863 0.0428-0.0428 $0.0956$ $0.1335$ 0.1827-0.1827 0.0603-0.0603 $0.0616$ $0.1813$ 0.1452-0.1452 0.0730-0.0730 $0.0187$ $0.1995$
GACC (TM) $0.3323$ 0.0635-0.0635 0.1405-0.1405 0.1283-0.1283 $0.3103$ 0.0250-0.0250 0.1426-0.1426 0.1427-0.1427 $0.2512$ $0.0486$ 0.1423-0.1423 0.1575-0.1575
GACC (RA) $0.1318$ 0.0065-0.0065 0.0592-0.0592 0.0661-0.0661 $0.1238$ $0.0132$ 0.0570-0.0570 0.0799-0.0799 $0.0975$ $0.0183$ 0.0329-0.0329 0.0830-0.0830
EV Owner 0.1684-0.1684 0.0350-0.0350 $0.0727$ $0.1307$ 0.1842-0.1842 0.0542-0.0542 $0.0459$ $0.1925$ 0.1580-0.1580 0.0817-0.0817 $0.0183$ $0.2215$
Age <50<50 0.0296-0.0296 0.0007-0.0007 $0.0131$ $0.0173$ 0.0363-0.0363 0.0034-0.0034 $0.0139$ $0.0258$ 0.0349-0.0349 0.0082-0.0082 $0.0102$ $0.0329$
Metropolitan 0.0568-0.0568 $0.0013$ $0.0250$ $0.0304$ 0.0511-0.0511 0.0026-0.0026 $0.0204$ $0.0333$ 0.0408-0.0408 0.0067-0.0067 $0.0129$ $0.0345$
Black $0.0665$ 0.0010-0.0010 0.0296-0.0296 0.0359-0.0359 $0.0462$ $0.0029$ 0.0182-0.0182 0.0309-0.0309 $0.0239$ $0.0039$ 0.0069-0.0069 0.0209-0.0209
Democrat 0.1832-0.1832 0.0084-0.0084 $0.0851$ $0.1066$ 0.1815-0.1815 0.0384-0.0384 $0.0798$ $0.1401$ 0.1393-0.1393 0.0660-0.0660 $0.0491$ $0.1562$
Similarly, the relationship between confidence about building charging infrastructure (measured in the 2023 survey) and purchase intent varies across the distribution of the latter. At the 80th percentile, confident respondents are 20.0 percentage points more likely to express strong purchase intent than those who are not confident. The corresponding differences are 18.1 percentage points at the median and 13.3 percentage points at the 20th percentile. Conversely, at the 80th percentile, those who are confident are 14.5 percentage points less likely to express no interest, with differences of 18.3 and 18.6 percentage points at the median and 20th percentile, respectively.
@@
Beliefs about EVs’ environmental benefits relative to gas-powered vehicles (measured in the 2021 survey) also show varying relationships with purchase intent across quantiles. Among respondents at the 80th percentile, those who believe EVs are environmentally better than gas-powered vehicles are 17.2 percentage points more likely to express strong purchase intent than those at the same percentile who do not share this belief. The corresponding differences are 13.8 percentage points at the median and 10.1 percentage points at the 20th percentile. Conversely, such respondents are 13.8 percentage points less likely to express no interest at the 80th percentile, compared to 17.7 and 19.6 percentage points less likely at the median and 20th percentile, respectively.
@@
Regarding views on government action, respondents who believe the government is doing too much or enough to reduce climate change are significantly less likely to express strong purchase interest, with this becoming more pronounced at higher percentiles. In the 2021 survey, those at the 80th percentile of purchase intent who hold this belief are 14.9 percentage points less likely to report strong purchase intent than those at the same percentile who do not. The gap is 14.1 and 10.7 percentage points at the median and 20th percentile, respectively. Likewise, at the 80th percentile, these respondents are 12.8 percentage points more likely to express no interest in EVs, with the gap widening to 19.7 and 21.9 percentage points at the median and 20th percentile, respectively. More recent surveys indicate even stronger covariate effects, with those who believe the government is doing too much approximately 33 and 25 percentage points more likely to express no interest in EVs at the 20th percentile and 80th percentile, respectively.
The influence of prior EV ownership also varies across quantiles. At the 80th percentile, respondents with prior EV ownership are 22.7 percentage points more likely to express strong purchase intent compared to those without prior ownership. This difference decreases to 20.8 and 16.2 percentage points at the median and 20th percentile, respectively. A similar pattern is observed in both the 2022 and 2023 surveys. Across surveys and quantiles, prior EV ownership is the strongest predictor for EV purchase intent.
@@
Finally, the relationship between political party affiliation and EV purchase intent varies considerably across quantiles. Results from the 2021 survey show that at the 80th percentile, Democrats are 14.3 percentage points more likely than Republicans to express strong purchase intent, with the difference narrowing to 9.6 percentage points at the 20th percentile. Similarly, at the 80th percentile Democrats are 8.3 percentage points less likely than Republicans to express no interest in EVs, but the difference is 12.4 percentage points at the 20th percentile. These patterns remain consistent in more recent surveys. The results indicate that political identity is closely linked to EV purchase intent across the entire distribution: at higher quantiles, Democrats are significantly more likely to express strong purchase intent, and at lower quantiles, they are far less likely to reject EVs outright.
Figure 3: The figure presents the covariate effects for some selected variables from the April 2021 and May-June 2023 surveys, spanning the 5th to 95th quantiles with 5th quantile increments. Legend: $\Delta\Pr(\textrm{Not at all likely})$:  ; $\Delta\Pr(\textrm{Not too likely})$:   ; $\Delta\Pr(\textrm{Somewhat likely})$:   ; $\Delta\Pr(\textrm{Very likely})$:  .
@@
To provide a clearer understanding of how covariate effects vary across quantiles, Figure 3 presents these for some selected variables from a series of ordinal quantile regression models, covering the 5th to 95th percentiles of EV purchase intent in 5-percentile increments. Each panel corresponds to a different covariate; with the solid gray, dashed gray, dashed black, and solid black lines representing the change in probability of reporting “not likely at all,” “not too likely,” “somewhat likely,” and “very likely,” respectively, for a given change in covariate across the quantiles. The figure clearly illustrates the substantial differences in covariate effects across the distribution.
@@
For example, the solid black line indicates that respondents at higher quantiles who have read or heard extensively about EVs, perceive them as environmentally beneficial, have owned one, identify as Democrats, or are confident in the US ability to develop infrastructure are much more likely to report strong purchase interest compared to those at lower quantiles. Conversely, the solid gray lines show that those who believe government climate efforts are excessive or adequate are significantly more likely to report no interest at lower quantiles and markedly less likely to report strong interest at higher quantiles.
5 Discussion
Our findings demonstrate how targeted strategies may be able to accelerate EV adoption. Exposure to EV information is associated with a 10.2 percentage point higher probability of reporting strong intent, with larger differences among respondents at the upper end of the intent distribution. It is also associated with an 8.5 point lower probability of reporting no intent. However, only 37.7% of respondents report high exposure, pointing to an opportunity for broader outreach through media, industry, and education. The smaller covariate effect of prior exposure at the 20th percentile suggests that simply raising awareness may not persuade consumers with low purchase interest. In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may influence purchase decisions.
Belief that EVs are better for the environment than gas-powered vehicles is associated with a 15.5 percentage point higher probability of reporting strong intent, with larger differences at the upper end of the distribution. It is also associated with a 19.5 point lower probability of reporting no interest, with larger differences at the lower end of the distribution. This pattern suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters and reduces skepticism among consumers with lower purchase interest. In the 2021 survey, 71.8% of respondents expressed this belief. Emphasizing environmental benefits—such as improved air quality—may resonate with both likely adopters and more skeptical respondents, broadening support for EV adoption.
Confidence in charging infrastructure is one of the strongest predictors of purchase intent. Respondents who express confidence that the US will build sufficient infrastructure are 19.1 percentage points more likely to report strong intent and 18.8 points less likely to report no interest. Quantile estimates suggest that confidence both strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent. Despite the importance of this factor, only 17.5% of respondents in the 2023 survey reported high confidence. Policymakers and industry can address this gap through continued investment in charging infrastructure-such as the National Electric Vehicle Infrastructure (NEVI) program-to improve affordability, accessibility, and reliability, as well as through effective communication of these efforts and of a long-term vision for the charging network.
Belief that the government is doing too much to address climate change is associated with a 14.8–19.7 point lower probability of strong intent and a 19.7–33.4 point higher probability of no interest. In the 2023 survey, 24.1% of respondents held this view. These findings suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonate across ideological lines. Rather than positioning EVs primarily as a climate solution, highlighting other benefits—such as long-term cost savings and air quality improvements—may engage a broader range of potential buyers. Our finding that Democrats are more likely to express stronger interest in EVs, further highlights the role of political identity and the value of emphasizing non-partisan benefits to broaden appeal. Moreover, for consumers wary of climate-focused policies, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns.
The results for other covariates point to further implications. Prior EV ownership is the strongest predictor of purchase intent, underscoring the importance of reducing barriers for first-time buyers. Potential options include subsidies, test-drive programs, and community-based education. Younger, urban, educated, and higher-income individuals report greater interest in EVs, suggesting that targeting information or advertising to such individuals may help raise adoption rates. By contrast, lower interest among women, US-born, and Black respondents may reflect cultural, structural, or informational barriers that could be addressed through inclusive policy design and outreach.